A cylindrical metal tube has a length of 50 cm and is open at both ends. Find the frequencies between 1000 Hz and 2000 Hz at which the air column in the tube can resonate. Speed of sound in air is 340 m s-1.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. According to the questions f base 1 first overtone of a closed organ pipe P base 1 = 3v/4l = 3* V/ 4 * 30 f base 2 fundamental frequency of a open organ pipe P base 2 = V/2l base 2 Here given 3V/4 * 30 = V/2l base 2 ⇒ l base 2 = 20 cm ∴length of the pipe P base 2 will be 20 cm